OpenStudy (anonymous):

Find the critical points of f(x)= x sqrt(x+3). I don't know how to simplify the derivative to find the critical point. Someone help please!

OpenStudy (abb0t):

Take the derivative of this function and find the roots.

OpenStudy (abb0t):

Meaning, find the zero's, and plug those values back into f(x).

OpenStudy (anonymous):

I did that but i get stuck with \[\sqrt{x+3} + \frac{ 1 }{ 2\sqrt{x+3}}\] I don't know how to simplify it.

OpenStudy (nincompoop):

erm ...

OpenStudy (turingtest):

you forgot an x in the numerator of the second term treat it like\[(x+3)^{1/2}+\frac12x(x+3)^{-1/2}\]which can be factored...

OpenStudy (turingtest):

\[(x+3)^{1/2}+\frac12x(x+3)^{-1/2}=(x+3)^{-1/2}(x+3+\frac x2)=0\]can you take it from here?

OpenStudy (anonymous):

I don't understand what you are doing in the 2nd part.

OpenStudy (anonymous):

Oh wait I understand it now! thank you for your help!

OpenStudy (nincompoop):

you're welcome

OpenStudy (turingtest):

Thanks nincompoop :p